OpenStudy (anonymous):

x^2 -2x-62=0 A. x=-1+-3(sqrt 7) B. x=1+-3(sqrt 7) C. x=2+_6(sqrt7) D. no real sol

OpenStudy (amorfide):

there are two ways either complete the square or use the quadratic formula completing the square \[x^{2} + bx + c=0\] to complete the square \[(x+\frac{ b }{ 2 })^{2}-(\frac{ b }{ 2 })^{2} + c=0\] in your case b=-2 c=-62

OpenStudy (amorfide):

rearrange to work out your answer

OpenStudy (anonymous):

I did that but I keep getting two answers and I neeed to pick one but I don't know which one @amorfide

OpenStudy (amorfide):

using the formula I gave you \[(x+\frac{ -2 }{ 2 })^{2} - (\frac{ -2 }{ 2 })^{2} - 62 = 0\] \[(x-1)^{2} - 1 - 62=0\] \[(x-1)^{2} -63=0\] \[(x-1)^{2}=63\]

OpenStudy (amorfide):

you can work out the rest yourself

OpenStudy (amorfide):

show me what you get and I will help